Kevin invested part of his 10,000 bonus in a certificate of deposit that paid 6% annual simple interest, and the remainder in a mutual fund the paid 11% annual simple interest. if his total interest for the year was 800, how much did kevin invest in the mutual fund?
Answer by jorel1380(3719)

x + y=10000
.06x + .11y=800
----------------
Multiply the second equation by 100:
6x+11y=80000
6x+6y =60000
--------------
5y=20000
y=4000
x=6000
-------------
Kevin invested $6000 in a CD, and $4000 in a mutual fund.